# A Novel Algorithm for Digital Lithological Mapping: Case Studies in Sri Lanka’s Mineral Exploration

### v1

@@ Abstract
Digital lithological mapping; Predictive modeling; Single-target identification; Site-specific signature; Pixel purity alignment
journal: Remote Sens. Environ.
@@ 1 Introduction
In recent years, spectral imaging (SI) has found applications in various domains, including climate and environmental monitoring (Henderson and Xia, (1997); Giuliani et al., (2017)), biodiversity studies (Randin et al., (2020)), ecosystem analysis (Donlon et al., (2012)), and food quality assessment (Bandara et al., (2020); Weerasooriya et al., (2020)). Airborne SI offers significant advantages for remote sensing (RS) due to its non-intrusive nature and its capacity to simultaneously capture data across a broad area at multiple spectral wavelengths. These characteristics underpin its utility in diverse applications.
@@
Prominent fields benefiting from RS include food and agriculture (Ekanayake et al., (2018)), ecology (Zhang et al., (2021)), hydrology (Chen et al., (2020)), and mineralogy (Grebby et al., (2014)). Remote sensing proves invaluable in these domains, demonstrating its versatility and applicability. Within mineralogy, RS plays a significant role in a process known as ’mineral indication’ (Yousefi et al., (2016)), which involves the identification of minerals. In this context, remote sensing primarily focuses on the spectral and radiometric properties of minerals, as opposed to their chemical and physical characteristics. Decades of research in mineral indication have produced spectral libraries used for mineral identification and classification, facilitating the generation of mineral distribution maps. The development in remote sensing (RS) techniques, as evidenced by research (Black et al., (2016);Feng et al., (2003); Grebby et al., (2011)), has opened up opportunities to enhance lithological mapping processes, thereby improving our ability to interpret compositions of geological sites.
@@
In this context, both hyperspectral images (HSIs) and multispectral images (MSIs) acquired through wide and narrow band sensors (del Rosario Gonzalez-Moradas and Viveen, (2020)) play vital roles in advancing digital lithological mapping (DLM). More specifically, within the field of Digital Lithological Mapping (DLM), Thermal Infrared (TIR) data obtained from strategically positioned multispectral sensors have proven to be exceptionally valuable. Notable examples include sensors like the Thermal Infrared Multispectral Scanner and the Advanced Space-borne Thermal Emission and Reflection Radiometer (ASTER), which have showcased the remarkable capability of TIR data in effectively discriminating a wide variety of minerals, with a particular emphasis on silicates.(Hubbard and Crowley, (2005); Jiang et al., (2013); Ninomiya et al., (2005)).
@@
Multispectral images (MSIs) are favored for DLM due to their inherent advantage of an improved signal-to-noise ratio (SNR) resulting from wide spectral band spacing. This spacing reduces cross-channel interference compared to hyperspectral images (HSIs). However, HSIs can be employed with algorithms that provide superior performance, even amidst noise interferences, owing to their spectral richness. It’s worth noting that the study sites in these investigations do not solely consist of minerals, devoid of other environmental features such as trees, soil, sand, etc. This factor makes it challenging to directly apply laboratory-generated spectral signatures (Ismail et al., (2014); Kruse and Perry, (2007); Pour et al., (2019)).
@@
Alternatively, spectral libraries (Rajendran et al., (2018)) can be employed to identify a group of known minerals. This is achieved by applying criteria such as maximum likelihood classification (Cabral et al., (2018)), spectral angle mapping (Dennison et al., (2004)), and spectral information divergence (Palsson et al., (2017)). These techniques prove effective because even if there is a minor error in identifying a mineral, it doesn’t lead to significant issues since the error is distributed across various mineral types. However, these same methods may not perform well when the goal is to pinpoint a specific mineral. This is due to the fact that when a mineral is mixed with other substances, its spectral characteristics undergo changes, rendering it distinct from what’s available in the library.
@@
In this study, a novel algorithm is introduced to generate the map of a target mineral of interest under the influence of various impurities. The algorithm consists of extracting a mineral representative signal from the HSIs, which will then be used to generate a map of the relative availability of the mineral in a given location. In contrast to the more traditional heuristic methods, one of the main steps in the algorithm is that of the autonomous separation of the mineral and the impurity signals from the HSI into the mineral and impurity representative subclasses. This subsequently opens up a way to utilize Fisher’s Discriminant Analysis, which maximizes the separability of these two subclasses. As a result of this maximization of separability, a quantity defined as relative availability of a specific target mineral in a given HSI pixel is calculated. Finally, by utilzing this quantity, a digital map of the mineral is generated.
@@
In order to firmly validate the performance of the algorithm, it was applied for three different minerals in four regions of Sri Lanka. The study considered Jaffna, Mannar, Pulmoddai and Murunkan, all of which are locations in the Northern parts of Sri Lanka, where mineral excavations are ongoing. The Jaffna area is popular for its Limestone, whereas Mannar and Pulmoddai both of which are well known for Ilmenite, and finally, Murunkan which is mostly famous for its clay like mineral, Montmorillonite. An extensive and rigorous study was conducted to compare the XRD and Magnetic Separation test results of the collected samples from these locations with the predicted availability of each site from the algorithm. The results demonstrate that the algorithm is capable of providing information with regards to the availability of a target mineral in a given location amidst impurities.
@@
  • 1.
    A novel hybrid machine-learning and signal-processing based algorithm for identifying and mapping any target mineral, utilizing remotely sensed hyperspectral imaging, regardless of the available mineral concentration.
  • 2.
    A case study which validate the results of the proposed algorithm, corroborating them with field surveys along with laboratory XRD and MS testing, is also presented. The case field survey data are collected spanning four sites (Mannar, Pulmoddai, Jaffna and Murunkan) and three minerals (Ilmenite, Limestone and Montmorillonite).
@@ 2 Related work
The introduction of hyperspectral sensors with multiple spectral channels has significantly improved the accuracy of remotely retrieved mineralogical and surface composition information (Niranjan et al., (2016); Swamy et al., (2017); and Zomer et al., (2009)). These sensors cover a wide spectral range, including the Thermal Infrared Region (TIR), and instruments like the Airborne Hyperspectral Scanner and others have been successful in mapping various minerals (Vaughan et al., (2003); Zhang et al., (2018)). Hyperion sensor data from the EO1 satellite has effectively mapped lithological units in India (Pal et al., (2020)). Pseudo-hyperspectral imagery, generated from Landsat data and regression-based models, has been used to identify metal deposit-related minerals (Hoang and Koike, (2016)). Hyperspectral images from satellites and airborne systems have also been employed to analyze minerals and rock microstructures while considering soil variability at different remote sensing scales (Van Ruitenbeek et al., (2019); Hecker et al., (2019)).
@@
In most digital lithological studies (Black et al., (2016); Ekanayake et al., (2019); Feng et al., (2003); Grebby et al., (2011); Tziolas et al., (2020)), the prevalent approach has been the utilization of generic laboratory-generated spectral libraries, with limited emphasis on generating site-specific signatures. Furthermore, when employing generic spectral signatures for lithological mapping, assumptions were often made regarding the presence of minerals at the site (Ninomiya and Fu, (2019); Yu et al., (2012)) or relied on pre-existing surveys of the site to confirm mineral composition (Pour et al., (2018); Xiong et al., (2011)). The former approach can be error-prone due to the lack of site-specificity in the signatures, while the latter serves primarily as a validation procedure, offering neither enrichment nor acceleration of the field survey (Rajan Girija and Mayappan, (2019)).
@@
The extraction of endmember signatures is often addressed and is a discrete research area in the literature for RS unmixing. These studies make use of unsupervised unmixing techniques in RS to derive the endmember signatures and the abundance map for the lithological map. For example, orthogonal subspace projection methods (Cheng et al., (2015); Li et al., (2015); Ren and Chang, (2000)) were used to classify spectral signatures following dimension reduction algorithms. Similarly, several dimensionality reduction-based methods such as PCA, ICA, and MNF have been applied in mineral exploration with ASTER satellite data (Shirmard et al., (2020)). Furthermore, accurate HSI classification has been performed via an improved version of the standard non-negative matrix factorization (NMF) algorithm incorporating fundamental notions of independence (Benachir et al., (2013); Sun et al., (2017)).
@@
Though existing techniques such as NMF-based unmixing (Rajabi and Ghassemian, (2013); Rathnayake et al., (2020); Wang et al., (2016)) or autoencoder architecture-based unmixing (Hua et al., (2020); Qi et al., (2020); Ranasinghe et al., (2020)) are superior at extracting the endmembers and estimating the corresponding abundances, these algorithms require knowledge about the number of endmembers to extract which was not available in the first place. Besides, for an algorithm to find feasible survey locations for a particular mineral through single-target identification, such information is superfluous. Similarly, the performance of blind source separation algorithms, such as the Pixel-Purity-Index salgorithm and Independent Component Analysis (ICA), which are used to extract sources of signals will be hindered due to the lack of pure mineral pixels and information on the endmembers. However, most of the study areas considered had a mineral composition that allowed accurate endmember extraction via automated unmixing algorithms. While the endmember extraction methods mentioned above are suitable for unpopulated geographical locations, in certain geographical regions, human intervention (Barbosa et al., (2003); Mendes et al., (2015); Vila Subiros et al., (2016)), has altered the composition of the soil with impurities. This alteration could make these methods unreliable in these areas.
@@
The availability of high-resolution remote sensing data has led to the accumulation of vast amounts of big data, opening new opportunities for data-driven discoveries. Currently, the integration of remotely sensed data with state-of-the-art data analytics, including machine learning, is significantly enhancing the field of geological mapping (Bachri et al., (2019)). For instance, certain remote sensing (RS) techniques and enhancement filters have been employed to trace geological structures and map hydrothermal alteration minerals using SPOT-5 and ASTER satellite data for Cu-Au prospecting (Ahmadirouhani et al., (2018)). Additionally, semi-automated support vector machines were utilized to map lithium (Li)-bearing pegmatites (Ahmadirouhani et al., (2018)), and the random forest classifier was applied for lithological classification based on remote sensing data (Bachri et al., (2019))
@@
According to the related works, DLM has been performed using spectral libraries of geographical regions that had a pure mineral composition, thereby the mineral signatures were more dominant than the spectral signatures of other non-mineral constituents or impurities. Although there is a limited number of studies to improve the SNR of remote sensing images, the detection of minerals under the strong influence of impurities is yet to be improved to the best of our knowledge. On the other hand, estimation of the abundances for mineral mapping has been performed under multi-target detection with unmixing techniques. But, the mineral map generation for a single mineral has not been under scrutiny since these unmixing techniques are susceptible to noise in the image.
@@
Addressing these gaps in the literature, this study introduces a comprehensive algorithm designed for the single-target detection of any mineral of interest in regions characterized by high impurity levels. The algorithm leverages Hyperspectral Images (HSIs) due to their superior spectral discrimination capabilities, despite their inherent challenges such as low SNR arising from cross-channel interference, as compared to Multispectral Images (MSIs) (Guo et al., (2023); Zhao et al., (2017)). Furthermore, the proposed algorithm makes use of existing hyperspectral libraries for single target mineral detection. Thus, this research addresses the identified deficiencies in the field, as outlined in Section , by contributing site-specific endmember generation techniques aimed at estimating the abundances of a specific mineral in the context of single-target detection.
@@ 3.1 Montmorillonite
Si_4O_10(OH)_2$\cdot$ nH_2O in which silicate layers sandwich a layer of aluminum oxide (Wanasinghe and Adikary, (2017)). The uses of Montmorillonite can be found in many different disciplines. It can significantly improve the needed performance when added to materials, polymers, and goods. It’s interesting to note how widely Montmorillonite is used in polymers and composites as a functional filler. (Herath, (1973)) Furthermore, it is used as a food additive for health and stamina, antibacterial activity, and as a sorbent for nonionic, anionic, and cationic dyes catalysts in organic synthesis, etc. In addition, it contributes as a plasticizer in sand casting, in drilling mud, and in electrical, heat, and acid-resistant porcelain. The prior research conducted on the distribution of different clay minerals in Sri Lanka provides evidence of Montmorillonite rich-clay deposits in the regions of Murunkan in Mannar, the vicinity around the Giant’s tank (Wanasinghe and Adikary, (2017))). The Murunkan region is characterized by its cracked earth, giving it its name, which means ”cracked earth” in Tamil. The name is derived from the black clay surface that tends to crack in hot, dry weather.
The clay deposits are abundant in the dry region and are concentrated around the water reservoirs which have been built as sources of water for cultivation. In fact, the prior geological research state that Montmorillonite is accumulated at the bottom of the reservoirs due to sedimentation (Herath, (1973)). This study focuses on the region of Giant’s Tank Sanctuary and Murugan in Mannar. The samples were collected from three sites; in the vicinity of the Giant tank, and the paddy fields on the roadsides of the Madawachchiya-Talaimannar highway.
@@ 3.2 Limestone
The East coast of the Jaffna Peninsula of Sri Lanka is such a geographical site where the lithological data of the dominant mineral is unavailable. The thick Jaffna Limestone is the dominant rock type that underlies the whole of the Jaffna Peninsula (Senaratne and Dissanayake, (1982)) and the surrounding islands. Formerly, Jaffna was an island composed of Miocene Limestone and the island was connected to the mainland by a spit formed of sediments brought by currents. These sediments were then carried to the eastern and northern coasts of the mainland and subsequently to the lagoon. The terrain of interest is a paleo-spit formed on the northeastern side of the peninsula which constitutes the shoreline, vegetation, and water bodies. However, the Limestone deposits are copious towards the West coast of Sri Lanka according to the findings of (Perera et al., (2020)). The study location selected in the work was in the vicinity of the Limestone region but lacks information about the existence of Limestone from previous studies.

### v2

@@ Abstract
Hyperspectral Imaging; Remote Sensing; Digital Lithological Mapping; Mineral exploration; Spectral signatures, Non-negative Least Squares
@@ 1 Introduction
In recent years, spectral imaging (SI) has found applications in various domains, including climate and environmental monitoring (Henderson and Xia 1997; Giuliani et al. 2017), biodiversity studies (Seto et al. 2004), ecosystem analysis (Yang 2005), and food quality assessment (Bandara et al. 2020; Weerasooriya et al. 2020). Airborne SI offers significant advantages for remote sensing (RS) due to its non-intrusive nature and its capacity to simultaneously capture data across a broad area at multiple spectral wavelengths. These characteristics underpin its utility in diverse applications.
@@
Prominent fields benefiting from RS include food and agriculture (Ekanayake et al. 2018), ecology (Zhang et al. 2021), hydrology (Chen et al. 2020), and mineralogy (Kusuma, Ramakrishnan, and Pandalai 2012). Remote sensing proves invaluable in these domains, demonstrating its versatility and applicability. Within mineralogy, RS plays a significant role in a process known as ’mineral indication’ (Yousefi et al. 2016), which involves the identification of minerals. In this context, remote sensing primarily focuses on the spectral and radiometric properties of minerals, as opposed to their chemical and physical characteristics. Decades of research in mineral indication have produced spectral libraries used for mineral identification and classification, facilitating the generation of mineral distribution maps. The development in remote sensing (RS) techniques, as evidenced by research (Black et al. 2016; Gersman et al. 2008; Grebby et al. 2011), has opened up opportunities to enhance lithological mapping processes, thereby improving our ability to interpret the compositions of geological sites.
@@
In this context, both hyperspectral images (HSIs) and multispectral images (MSIs) acquired through wide and narrow band sensors (del Rosario Gonzalez-Moradas and Viveen 2020) play vital roles in advancing digital lithological mapping (DLM). More specifically, within the field of Digital Lithological Mapping (DLM), Thermal Infrared (TIR) data obtained from strategically positioned multispectral sensors have proven to be exceptionally valuable. Notable examples include sensors like the Thermal Infrared Multispectral Scanner and the Advanced Space-borne Thermal Emission and Reflection Radiometer (ASTER), which have showcased the remarkable capability of TIR data in effectively discriminating a wide variety of minerals, with a particular emphasis on silicates.(Hubbard and Crowley 2005; Jiang, Lin, and Huang 2013; Bishop, Liu, and Mason 2011).
@@
Multispectral images (MSIs) are favored for DLM due to their inherent advantage of an improved signal-to-noise ratio (SNR) resulting from wide spectral band spacing. This spacing reduces cross-channel interference compared to hyperspectral images (HSIs). However, HSIs can be employed with algorithms that provide superior performance, even amidst noise interference, owing to their spectral richness. It’s worth noting that the study sites in these investigations do not solely consist of minerals, devoid of other environmental features such as trees, soil, sand, etc. This factor makes it challenging to directly apply laboratory-generated spectral signatures (Ismail et al. 2014; Kruse and Perry 2007; Pour et al. 2019).
@@
Alternatively, spectral libraries (Shanmugam and SrinivasaPerumal 2014) can be employed to identify a group of known minerals. This is achieved by applying criteria such as maximum likelihood classification (Cabral et al. 2018), spectral angle mapping (Dennison, Halligan, and Roberts 2004), and spectral information divergence (Palsson et al. 2017). These techniques prove effective because even if there is a minor error in identifying a mineral, it doesn’t lead to significant issues since the error is distributed across various mineral types. However, these same methods may not perform well when the goal is to pinpoint a specific mineral. This is due to the fact that when a mineral is mixed with other substances, its spectral characteristics undergo changes, rendering it distinct from what’s available in the library.
@@
In this study, a novel algorithm is introduced to generate the map of a target mineral of interest under the influence of various impurities. The algorithm consists of extracting a mineral representative signal from the HSIs, which will then be used to generate a map of the relative availability of the mineral in a given location. In contrast to the more traditional heuristic methods, one of the main steps in the algorithm is the autonomous separation of the mineral and the impurity signals from the HSI into the mineral and impurity representative subclasses. This subsequently opens up a way to utilise Fisher’s Discriminant Analysis, which maximises the separability of these two subclasses. As a result of this maximisation of separability, a quantity defined as the relative availability of a specific target mineral in a given HSI pixel is calculated. Finally, by utilising this quantity, a digital map of the mineral is generated.
@@
In order to firmly validate the performance of the algorithm, it was applied to three different minerals in four regions of Sri Lanka. The study considered Jaffna, Mannar, Pulmoddai, and Murunkan, all of which are locations in the Northern parts of Sri Lanka where mineral excavations are ongoing. The Jaffna area is popular for its Limestone, whereas Mannar and Pulmoddai both of which are well known for Ilmenite, and finally, Murunkan which is mostly famous for its clay like mineral, Montmorillonite. An extensive and rigorous study was conducted to compare the XRD and Magnetic Separation test results of the collected samples from these locations with the predicted availability of each site from the algorithm. The results demonstrate that the algorithm is capable of providing information with regards to the availability of a target mineral in a given location amidst impurities.
@@
  • A novel hybrid machine-learning and signal-processing based algorithm for identifying and mapping any target mineral, utilising remotely sensed hyperspectral imaging, regardless of the available mineral concentration.
  • A case study that validates the results of the proposed algorithm, corroborating them with field surveys along with laboratory XRD and MS testing, is also presented. The case field survey data were collected spanning four sites (Mannar, Pulmoddai, Jaffna and Murunkan) and three minerals (Ilmenite, Limestone and Montmorillonite).
@@ 2 Related work
The introduction of hyperspectral sensors with multiple spectral channels has significantly improved the accuracy of remotely retrieved mineralogical and surface composition information (Niranjan, Chaube, and Sarup 2016; Swamy, Asutkar, and Asutkar 2017; Zomer, Trabucco, and Ustin 2009). These sensors cover a wide spectral range, including the Thermal Infrared Region (TIR), and instruments like the Airborne Hyperspectral Scanner and others have been successful in mapping various minerals (Vaughan, Calvin, and Taranik 2003; Zhang et al. 2018). Hyperion sensor data from the EO1 satellite has effectively mapped lithological units in India (Pal, Rasmussen, and Porwal 2020). Pseudo-hyperspectral imagery, generated from Landsat data and regression-based models, has been used to identify metal deposit-related minerals (Hoang and Koike 2016). Hyperspectral images from satellites and airborne systems have also been employed to analyze minerals and rock microstructures while considering soil variability at different remote sensing scales (Van Ruitenbeek et al. 2019; Hecker et al. 2019).
@@
In most digital lithological studies (Black et al. 2016; Ekanayake et al. 2019; Pal et al. 2011; Grebby et al. 2011; Tziolas et al. 2020),the prevalent approach has been the utilization of generic laboratory-generated spectral libraries, with limited emphasis on generating site-specific signatures. Furthermore, when employing generic spectral signatures for lithological mapping, assumptions were often made regarding the presence of minerals at the site (Ninomiya and Fu 2019; Yu et al. 2012) or relied on pre-existing surveys of the site to confirm mineral composition (Pour et al. 2018; Xiong et al. 2011). The former approach can be error-prone due to the lack of site-specificity in the signatures, while the latter serves primarily as a validation procedure, offering neither enrichment nor acceleration of the field survey (Rajan Girija and Mayappan 2019).
@@
The extraction of endmember signatures is often addressed and is a discrete research area in the literature for RS unmixing. These studies make use of unsupervised unmixing techniques in RS to derive the endmember signatures and the abundance map for the lithological map. For example, orthogonal subspace projection methods (Cheng et al. 2015; Li, Chang, and Song 2015; Ren and Chang 2000) were used to classify spectral signatures following dimension reduction algorithms. Similarly, several dimensionality reduction-based methods such as PCA, ICA, and MNF have been applied in mineral exploration with ASTER satellite data (Shirmard et al. 2020). Furthermore, accurate HSI classification has been performed via an improved version of the standard non-negative matrix factorization (NMF) algorithm incorporating fundamental notions of independence (Benachir et al. 2013; Sun et al. 2017).
@@
Though existing techniques such as NMF-based unmixing (Rajabi and Ghassemian 2013; Rathnayake et al. 2020; Wang, Qian, and Tang 2016) or autoencoder architecture-based unmixing (Hua et al. 2020; Khajehrayeni and Ghassemian 2020; Ranasinghe et al. 2020) are superior at extracting the endmembers and estimating the corresponding abundances, these algorithms require knowledge about the number of endmembers to extract which was not available in the first place. Besides, for an algorithm to find feasible survey locations for a particular mineral through single-target identification, such information is superfluous. Similarly, the performance of blind source separation algorithms, such as the Pixel-Purity-Index salgorithm and Independent Component Analysis (ICA), which are used to extract sources of signals will be hindered due to the lack of pure mineral pixels and information on the endmembers. However, most of the study areas considered had a mineral composition that allowed accurate endmember extraction via automated unmixing algorithms. While the endmember extraction methods mentioned above are suitable for unpopulated geographical locations, in certain geographical regions, human intervention (Barbosa et al. 2003; Mendes et al. 2015; Vila Subiros et al. 2016) has altered the composition of the soil with impurities. This alteration could make these methods unreliable in these areas.
@@
The availability of high-resolution remote sensing data has led to the accumulation of vast amounts of big data, opening new opportunities for data-driven discoveries. Currently, the integration of remotely sensed data with state-of-the-art data analytics, including machine learning, is significantly enhancing the field of geological mapping (Bachri et al. 2019). For instance, certain remote sensing (RS) techniques and enhancement filters have been employed to trace geological structures and map hydrothermal alteration minerals using SPOT-5 and ASTER satellite data for Cu-Au prospecting (Ahmadirouhani et al. 2018). Additionally, semi-automated support vector machines were utilised to map lithium (Li)-bearing pegmatites (Ahmadirouhani et al. 2018), and the random forest classifier was applied for lithological classification based on remote sensing data (Bachri et al. 2019).
@@
According to the related works, DLM has been performed using spectral libraries of geographical regions that had a pure mineral composition, thereby the mineral signatures were more dominant than the spectral signatures of other non-mineral constituents or impurities. Although there is a limited number of studies to improve the SNR of remote sensing images, the detection of minerals under the strong influence of impurities is yet to be improved to the best of our knowledge. On the other hand, estimation of the abundances for mineral mapping has been performed under multi-target detection with unmixing techniques. But the mineral map generation for a single mineral has not been under scrutiny since these unmixing techniques are susceptible to noise in the image.
@@
Addressing these gaps in the literature, this study introduces a comprehensive algorithm designed for the single-target detection of any mineral of interest in regions characterized by high impurity levels. The algorithm leverages Hyperspectral Images (HSIs) due to their superior spectral discrimination capabilities, despite their inherent challenges such as low SNR arising from cross-channel interference, as compared to Multispectral Images (MSIs) ((Guo et al. 2023; Zhao et al. 2017). Furthermore, the proposed algorithm makes use of existing hyperspectral libraries for single-target mineral detection. Thus, this research addresses the identified deficiencies in the field, as outlined in Section , by contributing site-specific endmember generation techniques aimed at estimating the abundances of a specific mineral in the context of single-target detection.
@@ 3.1 Montmorillonite
The prior research conducted on the distribution of different clay minerals in Sri Lanka provides evidence of Montmorillonite rich-clay deposits in the regions of Murunkan in Mannar, the vicinity around the Giant’s tank (Wanasinghe and Adikary 2017). The Murunkan region is characterized by its cracked earth, giving it its name, which means ”cracked earth” in Tamil. The name is derived from the black clay surface that tends to crack in hot, dry weather.
The clay deposits are abundant in the dry region and are concentrated around the water reservoirs which have been built as sources of water for cultivation. In fact, the prior geological research state that Montmorillonite is accumulated at the bottom of the reservoirs due to sedimentation (Herath 1973). This study focuses on the region of Giant’s Tank Sanctuary and Murugan in Mannar. The samples were collected from three sites; in the vicinity of the Giant tank, and the paddy fields on the roadsides of the Madawachchiya-Talaimannar highway.
@@ 3.2 Limestone
The East coast of the Jaffna Peninsula of Sri Lanka is such a geographical site where the lithological data of the dominant mineral is unavailable. The thick Jaffna Limestone is the dominant rock type that underlies the whole of the Jaffna Peninsula (Senaratne and Dissanayake 1982) and the surrounding islands. Formerly, Jaffna was an island composed of Miocene Limestone and the island was connected to the mainland by a spit formed of sediments brought by currents. These sediments were then carried to the eastern and northern coasts of the mainland and subsequently to the lagoon. The terrain of interest is a paleo-spit formed on the northeastern side of the peninsula which constitutes the shoreline, vegetation, and water bodies. However, the Limestone deposits are copious towards the West coast of Sri Lanka according to the findings of (Perera et al. 2020). The study location selected in the work was in the vicinity of the Limestone region but lacks information about the existence of Limestone from previous studies.